. From pH of the following solutions, calculate the concentration (molarity) of [H+] and identify the solution as acidic, basic or neutral/nearly neutral (3 pts).

3.1) pH = 2.6

3.2) pH = 6.87
Answer #1

1)PH = -log [H+]

Given,ph =2.6

[H+] =10-ph

[H+]= 10-2.6

[H+] = 2.51 ×10-3

solution is acidic .Because ph is less than 7 .

2)PH = -log [H+] ,Given ph =6.87

[H+] =10-6.87

[H+]=1.349 ×10-7

Solution is nearly neutral because its ph is approximately near to 7.

Acid solution below 7 ph

Neutral solution 7 ph

Basic solution above 7 ph

Nearly neutral solution near to 7 ph